OpenStudy (anonymous):

What is the square root of -15 times the square root of -15?

OpenStudy (kinggeorge):

LATeX isn't working :( well, sqrt(-15) times sqrt(-15)=-15.

OpenStudy (anonymous):

OK, I thought there had to be and (i) in there to represent the - sqrt. thank you

OpenStudy (anonymous):

they cancel out im (-15^2) cancels out the parenthesis

OpenStudy (anonymous):

leavin you 15

OpenStudy (kinggeorge):

If you've gone over complex numbers before, it can be written as follows: sqrt(-15)sqrt(-15)= i^2 sqrt(15)^2=-1*15=-15

OpenStudy (anonymous):

OK Thx

OpenStudy (kinggeorge):

You're welcome.